### v1

# Looking for Tiny Defects via Forward–Backward Feature Transfer

Abstract
Motivated by efficiency requirements, most anomaly detection and segmentation (AD&S) methods focus on processing low-resolution images, e.g., $224\times 224$ pixels, obtained by downsampling the original input images. In this setting, downsampling is typically applied also to the provided ground-truth defect masks. Yet, as numerous industrial applications demand identification of both large and tiny defects, the above-described protocol may fall short in providing a realistic picture of the actual performance attainable by current methods. Hence, in this work, we introduce a novel benchmark that evaluates methods on the original, high-resolution image and ground-truth masks, focusing on segmentation performance as a function of the size of anomalies. Our benchmark includes a metric that captures robustness with respect to defect size, i.e., the ability of a method to preserve good localization from large anomalies to tiny ones. Furthermore, we introduce an AD&S approach based on a novel Teacher–Student paradigm which relies on two shallow MLPs (the Students) that learn to transfer patch features across the layers of a frozen vision transformer (the Teacher). By means of our benchmark, we evaluate our proposal and other recent AD&S methods on high-resolution inputs containing large and tiny defects. Our proposal features the highest robustness to defect size, runs at the fastest speed, yields state-of-the-art performance on the MVTec AD dataset and state-of-the-art segmentation performance on the VisA dataset.
1 Introduction
Image-based Anomaly Detection and Segmentation (AD&S) aims to identify anomalous samples and localize their defects. This task is particularly challenging in industrial applications where anomalies are varied and unpredictable, and nominal samples may be scarce. In these settings, AD&S is usually tackled in a cold-start fashion: the training procedure is unsupervised, with the train set comprising only images of nominal samples. Modern approaches for AD&S [25, 17, 10, 27, 8, 14, 29, 1] create a model of the nominal samples during training, while at inference time each test sample is compared to this nominal model and any discrepancy is interpreted as an anomaly. To reduce both training and inference time, and limit memory occupancy, most of these AD&S solutions create models based on low-resolution images obtained by downsampling the original input images. However, this approach is detrimental to the task since tiny anomalies could be lost due to strong downsampling, as may be observed in Fig. 1. Moreover, it is common practice to downsample also the ground-truth defect masks provided with the benchmarks. Accordingly, as shown in Fig. 1, the areas of defects get smaller, and tiny anomalies may even disappear from the ground-truth. Yet, since many industrial applications require the detection of both large and tiny defects, the practice described above might not accurately reflect the ability of current methods to localize defects of all sizes. Recently, EfficientAD [1] proposed a benchmark in which all the considered methods’ outputs are upsampled to the original ground-truth resolution, although all methods still process a low-resolution input.
In this work, we introduce a novel benchmark, referred to as Tiny Anomalies where all methods are evaluated on the original ground-truth resolution and, unlike [1], process high-resolution input images. Our benchmark focuses on evaluating segmentation performance as a function of the size of the anomalies and includes a novel metric that captures robustness with respect to the defect size, i.e., the ability to preserve localization performance from large anomalies to tiny ones.
Figure 1: Effects of downsampling on VisA. Tiny defects are no longer visible in both RGB and GT.
Furthermore, we propose a novel unsupervised AD&S approach that can process high-resolution input images at an exceedingly fast speed. Our approach relies on a frozen transformer backbone [22] and a novel Teacher–Student paradigm whereby lightweight MLPs (i.e., the Students) shared across all patch embeddings learn to mimic the transformation occurring between the layers of the transformer backbone (i.e., the Teacher) by observing only nominal samples. At inference time, for each patch, the actual features computed by the teacher are compared to those predicted by the students, with the discrepancies between the former and the latter highlighting the presence of anomalies.
Figure 2: Comparison between AD&S methods by the Tiny Anomalies benchmark. The metrics reported in the charts are described in Sec. 4. Values are normalized for better readability.
The strengths of our method come from several reasons. As far as the representation of small defects is concerned, a transformer backbone can effectively handle high-resolution inputs because, although it processes images by dividing them into patches, which results in a lower spatial size, the input information is not compressed, on the contrary, each patch is expanded to a higher dimensionality related to the internal representation of the transformer (e.g. RGB patches of $14\times 14\times 3$ pixels are mapped into $768$-dimensional embeddings in [22]). These features undergo subsequent transformations, through the different layers of the transformer, which further enrich the representation by increasing their level of contextualization. In particular, we rely on DINO-v2 [22] which has been trained on images of varying resolutions. For what concerns computational requirements, each of these rich, highly contextualized feature vectors can be processed independently through shared shallow MLPs which permits extremely fast batched processing.
As depicted in Fig. 2, our approach achieves the best segmentation performance on all defects, including tiny ones, demonstrates superior robustness, and is significantly faster than other methods. Our contribution can be summarized as follows: (i) we propose an evaluation protocol to establish a common ground, enabling comparison of performance across different AD&S methods; (ii) we introduce a novel benchmark to assess how effectively AD&S methods handle tiny anomalies; (iii) we propose a novel AD&S method that exhibits state-of-the-art performance on MVTec AD and state-of-the-art segmentation performance on VisA, while running at a remarkably higher speed than all competitors; (iv) our proposal achives state-of-the-art segmentation performance on a challenging few-shot AD benchmark built upon the VisA dataset.
2 Related work
Anomaly detection benchmarks.
During the last few years, several AD&S datasets have been released. The introduction of MVTec AD [4] kicked off the development of AD&S approaches for industrial applications. This dataset contains several industrial inspection scenarios, each comprising both train and test sets. Each train set contains only nominal images, while the test sets also contain anomalous samples. Such scenario represents realistic real-world applications where types and possible locations of defects are unknown during the development of AD&S algorithms. Later, the work was extended with the MVTec 3D-AD [6] dataset, which follows the same structure of MVTec AD, but also provides the pixel-aligned point clouds of the samples to address the AD&S in a multimodal fashion. Shortly afterwards the Eyecandies [7] dataset was released, which mimics the structure of MVTec 3D-AD by introducing a multimodal synthetic dataset containing images, point clouds and normals for each sample. To provide a more challenging scenario the VisA dataset [34] has been introduced, in which high-resolution images of complex scenes that can also contain multiple instances of the same object has been released. In the end, more task-specific datasets such as MAD [33] and MVTec LOCO [3] have been released. In particular, MAD [33] introduced a multi-pose dataset with images from different viewpoints covering a wide range of poses to tackle a pose-agnostic AD&S. MVTec LOCO [3] not only contains structural anomalies, such as dents or holes but also logical anomalies, which violations of logical constraints can be for instance a wrong ordering or a wrong combination of normal objects.
Anomaly detection solutions.
Several approaches have been proposed in the literature to perform AD&S. These solutions can be categorized based on the approach followed to model nominal samples. Normalizing Flows [23, 32, 26, 17, 10] based methods construct complex distributions by transforming a probability density via a series of invertible mappings. In particular, these methods extract features of normal images from a pre-trained model and transform the feature distribution into a Gaussian distribution during the training phase. At test time, after passing the extracted features through the Normalizing Flow, the features of abnormal images will deviate from the Gaussian distribution of the training phase, suggesting an anomaly. Lately, several solutions [25, 11, 2] that employ Memory Banks have been introduced. This category of solutions exploits well-known feature extractors trained on a large plethora of data [9, 22, 18] to model nominal samples. More in detail, during training, the feature extractor is kept frozen and used to compute features for nominal samples which are then stored in a memory bank. At test time, the features extracted from an input image are compared to those in the bank in order to identify anomalies. Despite their remarkable performance, these approaches suffer from slow inference speed, since each feature vector extracted from the input image needs to be compared against all the nominal feature vectors stored in the memory bank. Methods close to our solution which follows a Teacher–Student strategy [5, 31, 8, 28, 14, 1, 29] have also been proposed. In this family of solutions, the training phase involves a teacher model that extracts features from nominal samples and distils this knowledge to the student model, which learns to mimic the teacher’s feature extraction process. During the testing phase, differences between the features generated by the teacher model and those produced by the student model reveal the presence of anomalies. Recently, a multimodal approach [12] investigated the idea of mapping features from one modality to the other on nominal samples and then detecting anomalies by pinpointing inconsistencies between observed and mapped features. This solution leverages MLPs to learn a mapping between features coming from two different modalities, RGB images and point clouds. Differently, our solution does not necessarily require two modalities but can pinpoint anomalies by looking for discrepancies between extracted and transferred features of the same modality.
3 Method
As outlined in  Fig. 3, our approach follows a Teacher–Student paradigm in which the Teacher, $\mathcal{T}$, is represented by a frozen transformer encoder [22] while we define two students, referred to as Forward and Backward Transfer Network ($\mathcal{S}_{F}$ and $\mathcal{S}_{B}$) realized as shallow MLPs.
The Students are trained on nominal samples and learn to mimic the transformations between the patch embeddings occurring within the layers of the Teacher. In particular, the Forward Transfer Network learns to predict the patch embeddings computed by a layer of the transformer ($k$ in  Fig. 3), given the corresponding embeddings computed by a previous layer ($j$ in Fig. 3). Conversely, the Backward Transfer Network learns to predict the features calculated by the transformer at layer $j$ given the corresponding ones at layer $k$. The Student networks $\mathcal{S}_{F}$, $\mathcal{S}_{B}$ are shared across patch embeddings, i.e., both take as input the features associated with the patch $(i)$ at a layer $f_{j}^{(i)}$, $f_{k}^{(i)}$ and predict the corresponding features at the other layer $\hat{f}_{k}^{(i)}$, $\hat{f}_{j}^{(i)}$. At inference time, for all patch embeddings of the given test sample, the features predicted by the Students are compared to the ones extracted by the Teacher, with the discrepancies between the former and the latter providing the signal to highlight anomalies. In particular, as shown in  Fig. 3, the difference between the outputs from $\mathcal{S}_{F}$, $\mathcal{S}_{B}$ and the actual patch embeddings from layers $k$,$j$ of $\mathcal{T}$ yield two anomaly maps, $\Delta_{j}$ and $\Delta_{k}$, that are fused to obtain the final one. Due to the key novelty of our AD&S approach relying on the novel pretext task realized through $\mathcal{S}_{F}$, $\mathcal{S}_{B}$, we dub it Forward Backward Feature Transfer (FBFT).
The intuition behind our approach relies on the observation that as patch embeddings travel from shallower to deeper layers of a transformer encoder they become more and more contextualized, i.e., deeper representations capture more global information that helps singling out a patch based on the specific context provided by the input image. This complex contextualization function is performed progressively through the layers, with the different attention heads specializing in performing different tasks [30]. Hence, our Forward and Backward Transfer networks are trained to contextualize and de-contextualize patch embeddings according to the function, which we assume to be invertible, executed by the transformer between a pair of chosen layers. We conjecture that feature contextualization and de-contextualization are complex functions that do not admit a trivial solution, such as, e.g., the identity. Therefore, small-capacity networks trained only on nominal samples are unlikely to learn general functions that can yield correct predictions on out-of-distribution data, i.e. features extracted from anomalous patches. In the following, we provide additional details on our AD&S pipeline.
Figure 3: Proposed method. Given an RGB Image $I$, a frozen pre-trained transformer backbone $\mathcal{T}$ is leveraged to extract two sets of patch-aligned features $F_{j},F_{k}$, from different layers, one from a lower contextualization layer $j$ and one from a higher contextualization layer $k$. Then, a pair of feature transfer networks, $\mathcal{S}_{F},\mathcal{S}_{B}$, predict the extracted features from one layer to the other, processing the features at each patch independently. Lastly, extracted and transferred features are compared through a Euclidean distance, to create contextualization-specific anomaly maps, $\Delta_{j},\Delta_{k}$, that are then combined to obtain the final anomaly map $\Delta$.
Teacher.
As a first step, we provide as input to the teacher $\mathcal{T}$ an image $I$ with dimensions $H\times W\times C$, where $H$, $W$, and $C$ correspond to the height, width, and number of channels. In our framework, we employ a Transformer-based backbone that provides after each layer a set of features, one for each input patch processed by the backbone. Each feature, $f^{(i)}\in\mathbb{R}^{D}$, has dimension $D$ according to the inner representation employed by the backbone, while the number of features is $N=HW/P^{2}$, where the patch size is $P\times P$ pixels. During the forward pass, we extract two sets of features, $F_{j}=\bigl{\{}f^{(i)}_{j},i=1\cdots N\bigr{\}}$ and $F_{k}=\bigl{\{}f^{(i)}_{k},i=1\cdots N\bigr{\}}$, from two different layers of the backbone, i.e. layers $j$ and $k$, with $j<k$.
Students.
The two sets of features extracted by the Teacher are processed by a pair of Forward and Backward Transfer networks, $\mathcal{S}_{F}$ and $\mathcal{S}_{B}$, which in our architecture represent the students. $\mathcal{S}_{F}$ maps a feature vector from a less contextualized layer $j$ to a more contextualized layer $k$, while $\mathcal{S}_{B}$ does the opposite. Each network predicts the features of one layer from the corresponding ones extracted from the other, processing each patch location independently. Thus, given a patch location $(i)$ and the corresponding features $f^{(i)}_{j}$ and $f^{(i)}_{k}$, the features predicted by the students can be expressed as:
$\hat{f}^{(i)}_{k}=\mathcal{S}_{F}(f^{(i)}_{j})\quad\hat{f}^{(i)}_{j}=\mathcal{S}_{B}(f^{(i)}_{k})$ (1)
where $\mathcal{S}_{F}$ and $\mathcal{S}_{B}$ are parametrized as MLPs, shared across all patches. By processing all patches, we obtain the two sets of transferred features: $\hat{F}_{j}=\bigl{\{}\hat{f}^{(i)}_{j},i=1\cdots N\bigr{\}}$ and $\hat{F}_{k}=\bigl{\{}\hat{f}^{(i)}_{k},i=1\cdots N\bigr{\}}$.
Training.
During training, the weights of $\mathcal{S}_{F}$ and $\mathcal{S}_{B}$ are optimized considering only images of nominal samples of a specific class from a dataset. As loss function, for both networks we employ the cosine distance between the features extracted from the backbone at the considered layers and the transferred ones. Thus, the per-patch losses are:
$\mathcal{L}^{(i)}_{j}\Bigl{(}f_{j}^{(i)},\hat{f}_{j}^{(i)}\Bigr{)}=1-\frac{f_{j}^{(i)}\cdot\hat{f}_{j}^{(i)}}{\|f_{j}^{(i)}\|\|\hat{f}_{j}^{(i)}\|}\quad\mathcal{L}^{(i)}_{k}\Bigl{(}f_{k}^{(i)},\hat{f}_{k}^{(i)}\Bigr{)}=1-\frac{f_{k}^{(i)}\cdot\hat{f}_{k}^{(i)}}{\|f_{k}^{(i)}\|\|\hat{f}_{k}^{(i)}\|}$ (2)
Inference.
At inference time, the image under analysis is processed by the transformer backbone and the features extracted from the two layers, $F_{j}$ and $F_{k}$ are provided as input to the Forward and Backward Transfer networks to obtain the corresponding transferred features, $\hat{F}_{j}$ and $\hat{F}_{k}$. The Euclidean distance is then employed to compute the patch-wise differences between extracted and transferred features $\Delta^{(i)}_{j},\Delta^{(i)}_{k}$:
$\Delta^{(i)}_{j}=\|f^{(i)}_{j}-\hat{f}^{(i)}_{j}\|_{2}\quad\Delta^{(i)}_{k}=\|f^{(i)}_{k}-\hat{f}^{(i)}_{k}\|_{2},\quad i=1\ldots N$ (3)
which are then fused by multiplying those corresponding to the same patch:
$\Delta^{(i)}=\Delta^{(i)}_{j}\cdot\Delta^{(i)}_{k},\quad i=1\ldots N$ (4)
The set of fused differences, $\Delta^{(i)}$, is reshaped as a $\sqrt{N}\times\sqrt{N}$ anomaly map according to the positions of the patches within the input image. This map is then up-sampled to $H\times W$, i.e. the spatial size of the input image, by bilinear interpolation and successively smoothed according to common practice [25, 12, 29, 21]. The global anomaly score required to perform sample-level anomaly detection is computed as the mean value of the top $M$ values of the final anomaly map $\Delta$.
4 Tiny Anomalies benchmark
Most of the existing AD&S proposals operate on downsampled versions of the images, mainly due to computational constraints. However, as highlighted in Sec. 1 this practice is detrimental for tiny defects which tend to be reduced or in the worst case disappear in the downsampled versions of the input image and the ground truth. To this end, we introduce a cumulative quartile-based benchmark designed to analyze AD&S performance across defects of varying sizes.
Datasets.
To define such a benchmark we rely on two AD&S datasets: VisA [34] and MVTec AD [4]. The VisA [34] dataset provides images of varying resolution, with the height spanning from $1284$ to $1562$ pixels and anomalies as tiny as $1$ pixel and up to $478781$ pixels.
Figure 4: Anomaly size distribution.
The dataset contains $10821$ images of $12$ objects across $3$ domains, with challenging scenarios including complex structures in objects, multiple instances, and pose variations. Between the provided images, $9621$ are nominals while $1200$ contains defects. The MVTec AD dataset mimics real-world industrial inspection scenarios and includes $5354$ images, with heights spanning from $700$ to $1024$ pixels and anomalies ranging from $24$ pixels to $517163$ pixels. The images pertain to $15$ objects exhibiting $73$ different types of anomalies for $1888$ anomalous samples.
Both VisA and MVTec AD provide pixel-accurate ground truths for each anomalous sample to allow us to evaluate methods on both anomaly classification and segmentation. As highlighted in Fig. 4, VisA features a significantly wider range of anomaly sizes and includes tiny defects. As a result, downsampling the ground-truths to $224\times 224$ pixels, i.e., the most commonly employed inference and evaluation size in present literature, yields a reduction in the number of defects of 21.42% and 0.37% for VisA and MVTec AD, respectively.
4.1 Metrics.
The metrics defined in our benchmark build upon the evaluation proposed by MVTec AD [4] and VisA [34]. These two datasets, assess the image anomaly detection performance employing the Area Under the Receiver Operator Curve (I-AUROC) computed on the global anomaly score. As for segmentation performance, the Area Under the Per-Region Overlap (AUPRO) on the anomaly map is computed, with the integration threshold set to $0.3$. Recently, [1, 12] have proposed to compute the AUPRO considering a tighter threshold, i.e., $0.05$. We will consider both metrics and denote AUPROs with integration thresholds $0.3$ and $0.05$ as AUPRO@30%, and AUPRO@5%, respectively.
Performance across defects sizes.
To highlight the capability of each method to segment defects with varying sizes, in our novel benchmark, we introduce a variation of the AUPRO metric. In particular, for each object in a dataset, we compute the anomaly size distribution and partition it in cumulative quartiles, denoted as $Q_{1},Q_{2},Q_{3},Q_{4}$. These cumulative quartiles are associated with sets that contain only anomalies with a size smaller than or equal to the considered quartile. Hence, the set associated with Q4 consists of all anomalies, while Q1 includes only the smallest ones. Then, we calculate the AUPRO@30% and AUPRO@5% on each set, with the segmentation metrics associated with Q4 being the already described segmentation metrics adopted in current benchmarks.
Robustness.
We introduce a novel metric, $\rho$, aimed at assessing the robustness of a method w.r.t. the size of the defects in a dataset. In particular, $\rho$ captures the ability of a method to segment accurately tiny defects as well as larger ones. Accordingly, we define the robustness as:
$\rho=w\cdot\left(1-s\right),\quad s=\frac{|\text{AUPRO}(Q_{4})-\text{AUPRO}(Q_{1})|}{\max\bigl{(}\text{AUPRO}(Q_{1}),\text{AUPRO}(Q_{4})\bigr{)}},\quad w=\frac{1}{4}\cdot\sum_{i=1}^{4}\text{AUPRO}(Q_{i})$ (5)
[…]
4.2 Evaluation protocol
We evaluate our proposal, FBFT, alongside with several state-of-the-art AD&S methods, such as PatchCore [25], SimpleNet [21], EfficientAD [1] and RD++ [29]. As our benchmark emphasizes the ability to deal with tiny anomalies, we provide as input to all methods images at the highest resolution that would enable execution on a single GPU, so as to avoid or minimize downsampling. In particular, we could handle input images up to $1036\times 1036$ pixels with EfficientAD, RD++ and FBFT, while the highest input resolution for PatchCore and SimpleNet was found to be $512\times 512$ pixels.
As described in [1], the results reported in SimpleNet [21] are obtained by repeatedly evaluating the metrics on all test images during training to select the best check-point. Analyzing the official implementation, we noticed how this protocol has been followed also by RD++ [29]. However, in real-world settings, the test data is not available at training time. Thus, to avoid overestimating the actual performance of the models, we disable the above check-point selection mechanism, train all methods for a fixed number of epochs and evaluate the model obtained at the last checkpoint. For [1, 21, 29], we train for the number of epochs specified in the official implementations.
PatchCore [25] employs a centre-crop of the input images since in MVTec AD, most of the defects lie within this cropped area. However, in a real-world scenario, anomalies can occur outside of this area, thus, we disable this strategy as it implies knowledge about the location of anomalies in the test set.
As anticipated in Sec. 1, we compute all metrics based on the original ground-truths provided with the datasets, which have the same resolution as the original input images. Hence, we do not downsample the ground-truths to the input image size processed by a method. If a method outputs an anomaly map at a smaller resolution than the original ground-truth, we bilinearly upsample the anomaly map to the same resolution as the ground-truth in order to calculate all metrics.
Some methods, including ours, must add padding to the input image in order to adapt it to the input size of the employed backbone. However, we remove these extra pixels from the final anomaly maps as, otherwise, they usually decrease the False Positive Rate (and thus artificially ameliorate the segmentation metrics) because they tend to yield very low anomaly scores. Finally, we calculate the AUPRO considering all the samples in the test set, both nominal and anomalous  [^1] We noticed that official code from [25], calculates the AUPRO only on anomalous test samples, obtaining higher scores since the false positive rate is inherently lower with this protocol..
5 Experiments
Implementation details.
EfficientAD [1] proposes two variants: EfficientAD-S and EfficientAD-M. We consider the latter since it provides better AD&S performance. To implement our Feature Extractor we employ DINO-v2 ViT-B/14 [22] pre-trained on a large, curated and diverse dataset of 142 million images, comprising ImageNet-22k [15, 24]. Thus, $\mathcal{T}$ processes $1036\times 1036\times 3$ RGB images and outputs $74\times 74\times 768$ feature maps. Both $\mathcal{S}_{F}$ and $\mathcal{S}_{B}$ consist of three linear layers, each but the last one followed by GeLU activations. The number of units per layer is $768$ for both $\mathcal{S}_{F}$ and $\mathcal{S}_{B}$. The two networks are trained jointly for $50$ epochs using Adam [20] with a learning rate of $0.001$. We select the layers $j=8$ and $k=12$ to realize the Feature Transfer Networks. A detailed ablation study on the choice of the best pair of layers is reported in the supplemental material. We employed $M=0.001\cdot H\cdot W$ to attain the number of pixels used to calculate the global anomaly score. We conducted all the experiments on a single NVIDIA GeForce RTX 4090.
Anomaly detection and segmentation.
Following the proposed evaluation protocol, described in Sec. 4, we evaluate our method on VisA and MVTec AD, reporting the results in Tab. 1. Our approach achieves the best segmentation results on the VisA dataset, with $0.952$ AUPRO@30% and $0.787$ AUPRO@5% and the state-of-the-art in both detection and segmentation on the MVTec AD dataset, with $0.988$ I-AUROC, $0.945$ AUPRO@30%, and $0.782$ AUPRO@5% Regarding detection performance on VisA, our method attains results comparable to the runner-up ($0.968$ of EfficientAD vs. $0.964$ of Ours). The supplemental material provides the detailed per-class metrics for each method.
Table 1: I-AUROC, AUPRO30@and AUPRO5@on VisA and MVTec AD for several AD&S methods. Average metrics of all classes on the respective test set. Best results in bold, runner-ups underlined. All methods are trained and tested at high resolution.
Algorithm VisA MVTec AD
I-AUROC AUPRO@30% AUPRO@5% I-AUROC AUPRO@30% AUPRO@5%
PatchCore [25] 0.982 0.752 0.542 0.983 0.937 0.701
SimpleNet [21] 0.904 0.718 0.469
EfficientAD [1] 0.968 0.937 0.777 0.965 0.920 0.757
RD++ [29] 0.930 0.907 0.758 0.915 0.901 0.716
FBFT (Ours) 0.964 0.952 0.787 0.988 0.945 0.782
In Fig. 5, we depict some qualitative results on the VisA dataset. Our method provides more localized anomaly scores compared to EfficientAD [1], i.e., the second-best method on VisA. For instance, by looking at the capsules example, our anomaly score peak is centered on the anomaly differently from [1]. Further qualitative results are reported in the supplemental material.
RGB GT EfficientAD FBFT RGB GT EfficientAD FBFT
candle
Refer to caption Refer to caption Refer to caption Refer to caption
cashew
Refer to caption Refer to caption Refer to caption Refer to caption
capsules
Refer to caption Refer to caption Refer to caption Refer to caption
chewinggum
Refer to caption Refer to caption Refer to caption Refer to caption
fryum
Refer to caption Refer to caption Refer to caption Refer to caption
macaroni1
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: VisA dataset qualitative results. All methods are trained and tested at high resolution.
Cumulative quartiles based anomaly segmentation.
We report in Tab. 2 the analysis on VisA and MVTec AD of the performance w.r.t. anomaly size using the cumulative quartile metrics defined in Sec. 4. The results highlight that the defect size impacts the segmentation metrics, especially for the tiniest ones, i.e., the anomalies in $Q_{1}$. Nevertheless, our method is the best across all quartiles, with a notable gap compared to the second-best method on $Q_{1}$ on VisA, which is the dataset with the highest frequency of tiny defects (e.g., AUPRO30@$0.935$ Ours vs. $0.890$ EfficientAD). Additionally, our method is remarkably stable and robust across quartiles. For instance, on VisA, we go from $0.758$ to $0.730$ AUPRO5@%, losing only $2.6\%$ segmentation quality, remarkably less than the runner-up method, EfficientAD, which decreases its performance of $6.7\%$, from $0.743$ to $0.693$ AUPRO5@%.
Table 2: Quartile-based benchmark. Best results in bold, runner-ups underlined.
Algorithm Dataset AUPRO@30% AUPRO@5%
[…]
PatchCore [25] VisA 0.703 0.720 0.740 0.752 0.728 0.679 0.484 0.492 0.518 0.542 0.509 0.454
SimpleNet [21] 0.658 0.668 0.696 0.718 0.685 0.627 0.390 0.399 0.435 0.469 0.423 0.351
EfficientAD [1] 0.890 0.923 0.933 0.937 0.920 0.873 0.693 0.741 0.763 0.777 0.743 0.662
RD++ [29] 0.867 0.898 0.906 0.907 0.894 0.853 0.710 0.740 0.755 0.758 0.740 0.692
FBFT (Ours) 0.935 0.941 0.946 0.952 0.943 0.926 0.730 0.749 0.768 0.787 0.758 0.702
PatchCore [25] MVTec AD 0.924 0.932 0.935 0.937 0.932 0.918 0.653 0.677 0.691 0.701 0.680 0.633
EfficientAD [1] 0.922 0.925 0.925 0.920 0.923 0.920 0.758 0.769 0.767 0.757 0.762 0.760
RD++ [29] 0.946 0.922 0.918 0.901 0.921 0.952 0.782 0.752 0.744 0.716 0.748 0.684
FBFT (Ours) 0.958 0.948 0.947 0.945 0.949 0.986 0.806 0.798 0.795 0.782 0.795 0.783
Inference time.
We report in Tab. 3 the inference time and main AD&S metrics on VisA for our method and state-of-the-art approaches [25, 21, 1, 29]. Using the same machine, we compute the speed in ms per sample as the average across all the test samples of the VisA dataset. For each method, we compute the inference time, from when the sample is available on the GPU to the computation of the anomaly scores, after a GPU warm-up, synchronizing all threads before estimating the total inference time. Our approach attains state-of-the-art anomaly segmentation performance, namely AUPRO@30%, $\rho$@30%, AUPRO@5%, and $\rho$@5%, while being extremely fast. We highlight that, even though PatchCore attains the best detection performance on the VisA dataset, it largely falls behind in terms of segmentation performance (AUPRO@30%=0.752 of PatchCore vs. AUPRO@30%=$0.952$ of Ours, AUPRO@5%=$0.542$ of PatchCore vs. AUPRO@5%=$0.787$ of Ours), and inference speed ($227.230$ ms of PatchCore vs. $1.786$ ms of Ours).
Table 3: Inference speed and AD performance on VisA dataset. Inference Time in ms per sample. Best results in bold, runner-ups underlined. All methods are trained and tested at high-resolution.
Algorithm Inference Time I-AUROC AUPRO@30% $\rho$@30AUPRO@5% $\rho$@5%
PatchCore [25] 227.230 0.982 0.752 0.679 0.542 0.454
SimpleNet [21] 560.617 0.896 0.718 0.627 0.469 0.351
EfficientAD [1] 82.367 0.968 0.937 0.873 0.777 0.662
RD++ [29] 63.176 0.930 0.907 0.853 0.758 0.692
FBFT (Ours) 1.786 0.964 0.952 0.926 0.787 0.702
Few-shot anomaly detection and segmentation.
As mentioned in Sec. 1, collecting many nominal samples in most industrial scenarios can be extremely expensive or unfeasible. Also, frequent production changeover requires fast adaptation. For these reasons, a beneficial property of AD&S methods is the ability to create a model of the nominal data even with few samples. We define a few-shot benchmark – based on the VisA dataset – randomly selecting $5$, $10$, and $50$ images from each category as training data. We train the compared methods [25, 21, 1, 29] along with our proposed approach on these samples, and we test them on the entire test set, with the evaluation protocol proposed in Sec. 4, reporting the results in Tab. 4. We obtain the best segmentation performance for both metrics (AUPRO@30% and AUPRO@5%) in all the few-shot settings, significantly improving the most challenging segmentation metrics (+$0.167$ AUPRO@5% on 5-shot) and retaining a stable segmentation performance (AUPRO@30% always above 0.9) across the various settings. These results confirm the ability of our method to optimize feature transfer networks even from a few nominal samples, thanks to the patch–independent processing enabled by the MLPs.
Table 4: Few-shot AD&S performance. Best results in bold, runner-ups underlined.
Algorithm Full 50-shot 10-shot 5-shot
I-AUROC AUPRO@30% AUPRO@5% I-AUROC AUPRO@30% AUPRO@5% I-AUROC AUPRO@30% AUPRO@5% I-AUROC AUPRO@30% AUPRO@5%
PatchCore [25] 0.982 0.752 0.542 0.959 0.724 0.485 0.948 0.704 0.459 0.916 0.698 0.455
SimpleNet [21] 0.896 0.718 0.469 0.917 0.758 0.430 0.883 0.725 0.408 0.862 0.691 0.377
EfficientAD [1] 0.968 0.937 0.777 0.831 0.854 0.569 0.816 0.806 0.469 0.810 0.834 0.511
RD++ [29] 0.930 0.907 0.758 0.776 0.861 0.563 0.615 0.733 0.303 0.555 0.654 0.253
FBFT (Ours) 0.964 0.952 0.787 0.927 0.934 0.743 0.897 0.910 0.695 0.879 0.901 0.678
Features visualization.
In Fig. 6, we show the contextualized feature maps before $f_{j},f_{k}$ and after $\hat{f}_{j},\hat{f}_{k}$ the feature transfer, as well as their $\Delta_{j},\Delta_{k}$ and final anomaly maps $\Delta$, for a nominal (top) and an anomalous (bottom) test sample of VisA. In the nominal case, we can notice how the features before and after the feature transfer look similar, resulting in low anomaly scores. In the anomalous case, as the extracted features $f_{j}$, $f_{k}$ fall out of the nominal distribution, the feature transfer network fails to contextualize or decontextualize them, resulting in erroneously reconstructed features $\hat{f_{j}}$ and $\hat{f_{k}}$. Thus, by analyzing the discrepancy between the original and reconstructed features, we produce accurate anomaly maps. Furthermore, after the combination, the overall anomaly map $\Delta$ exhibits less noise, thanks to the product-based aggregation employed in this work.
RGB $f_{j}$ $\hat{f}_{j}$ $\Delta_{j}$ $f_{k}$ $\hat{f}_{k}$ $\Delta_{k}$ $\Delta$
Nominal
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Anomalous
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Features visualization. Channels average of feature maps before and after feature mapping.
6 Discussion
We introduced a benchmark to evaluate the stability of existing methods in segmenting anomalies of different sizes, spanning from very tiny to larger ones, along with a fair and sound evaluation protocol to assess the performance. We hope to establish this protocol as a common practice when evaluating AD&S metrics, to standardize results and render methods truly comparable. In addition, we introduced a fast and lightweight approach based on forward–backward transfer networks that process features extracted from layers with different contextualization levels of a transformer backbone. The proposed solution achieves the best segmentation results on the VisA dataset, both on the classical benchmark and the proposed one, while also running at a remarkably faster speed, compared to existing AD&S approaches. Also, it exhibits state-of-the-art performance on the MVTec AD dataset. Lastly, our approach outperforms competitors in segmentation performance also when considering a more challenging few-shot scenario built upon the VisA dataset.
Though transformers can create a rich representation for each patch of the input high-resolution image, a limitation of our method resides in the small spatial size of the output anomaly map, which is constrained by the leveraged backbone and demand bilinear upsampling to the input resolution. An idea to enhance the introduced framework would be to employ strategies that can yield high-resolution feature maps, such as NaViT [13], which, by efficiently performing repeated inference across an image, can output higher-resolution feature maps, or FeatUp [16], which can restore lost spatial information in deep features, yielding feature maps with the same spatial size as the input.
Appendix A Supplemental material
In this supplemental material, we provide additional quantitative and qualitative results to validate the performance of the proposed approach.
A.1 Ablation on the layers considered for the forward and backward feature transfer networks.
We investigate the impact of transferring features from different levels of the transformer architecture, i.e., layers $j$ and $k$ described in Sec. 3. As shown in Tab. 5, transferring features between layers with high contextualization, i.e., the last four layers, begets better detection and segmentation results, with the transfer between $j=8$ and $k=12$ providing the best performance. We also observe that transferring features from closer layers, such as $j=11$ and $k=12$, can harm the performance. We believe that being the function learned by a single transformation layer smooth [19], the task of transferring between two close layers is simpler, thus, it might overgeneralize to anomalous samples, leading to worse performance. Conversely, between two farther layers, the function is highly non-linear. Nevertheless, the performance is quite stable after the layer $8$, independently of the employed layers.
Table 5: Layers Ablation. Best results in bold, runner-ups underlined.
Layers I-AUROC AUPRO@30% AUPRO@5%
$[1,4]$ 0.906 0.828 0.570
$[4,8]$ 0.940 0.941 0.773
$[8,12]$ 0.964 0.952 0.787
$[10,12]$ 0.960 0.950 0.784
$[11,12]$ 0.956 0.946 0.774
A.2 Ablation on the loss employed to optimize the forward and backward feature transfer networks.
Tab. 6 report the results obtained by the proposed framework considering different losses for the optimization of the forward and backward feature transfer networks, i.e., the Students MLPs. Even though minimizing the cosine distance achieves better performance it is worth noticing that the results are very similar despite the loss function employed.
Table 6: Loss ablation. Best results in bold, runner-ups underlined.
Loss I-AUROC AUPRO@30% AUPRO@5%
Cosine distance 0.964 0.952 0.787
L2 distance 0.954 0.950 0.786
A.3 Ablation on single contextualization-specific anomaly map.
We present in Tab. 7 the anomaly detection and segmentation results achieved by the proposed methods when considering the single anomaly separately. These results show that by combining anomaly maps from different layers of contextualization we are able to achieve better performance with respect to the single contextualization-specific anomaly maps. Indeed, the combination of multiple maps can be seen as an ensembling of different AD&S models.
Table 7: Anomaly map evaluation ablation. Best results in bold, runner-ups underlined.
Anomaly map I-AUROC AUPRO@30% AUPRO@5%
$\Delta_{k}\cdot\Delta_{j}$ 0.964 0.952 0.787
$\Delta_{j}$ 0.953 0.943 0.773
$\Delta_{k}$ 0.949 0.925 0.745
A.4 Ablation on different input resolution sizes during training.
We report in Tab. 8 the anomaly detection and segmentation performance achieved by the proposed methods when different input resolution sizes are considered for the feature extractor. The same resolution is used during training and inference, while the evaluation is performed by upsampling the output to the original full resolution of the ground-truth. From these results, it is possible to appreciate that the proposed solution is able to exploit the higher resolution and correctly detect and segment the majority of samples, with an I-AUROC of $0.964$ at full resolution, compared to $0.899$ when providing low-resolution images. The same trend can be observed for the localization metrics, i.e., AUPRO@30% and AUPRO@5%.
Table 8: Ablation on the input resolution employed at training time. Best results in bold, runner-ups underlined.
Training resolution I-AUROC AUPRO@30% AUPRO@5%
$224\times 224$ 0.899 0.830 0.562
$518\times 518$ 0.952 0.944 0.779
$1036\times 1036$ 0.964 0.952 0.787
A.5 Full results on VisA
For the sake of completeness, in Tab. 9 we report the per-class detection and segmentation performance, previously summarized in Tab. 1 of the main paper. Results are reported for our solution and the state-of-the-art methods on the VisA dataset.
Table 9: I-AUROC and AUPRO30@on the VisA dataset for several AD&S methods. Best results in bold, runner-ups underlined. All methods are trained and tested at high-resolution.
Algorithm candle capsules cashew chewinggum fryum macaroni1 macaroni2 pcb1 pcb2 pcb3 pcb4 pipe_fryum Mean
I-AUROC PatchCore [25] 0.986 0.937 0.990 0.991 0.993 0.997 0.934 0.980 0.988 0.996 0.998 0.998 0.982
SimpleNet [21] 0.964 0.769 0.972 0.984 0.922 0.809 0.618 0.984 0.956 0.949 0.937 0.986 0.904
EfficientAD [1] 1.000 0.884 0.933 0.996 0.957 0.947 0.967 0.991 0.971 0.972 1.000 1.000 0.968
RD++ [29] 0.846 0.935 0.862 0.838 0.966 0.964 0.897 0.935 0.972 0.980 0.982 0.990 0.930
FBFT (Ours) 0.958 0.992 0.972 0.996 0.988 0.931 0.885 0.980 0.938 0.956 0.976 0.999 0.964
AUPRO@30% PatchCore [25] 0.955 0.575 0.912 0.670 0.836 0.349 0.340 0.941 0.864 0.703 0.910 0.969 0.752
SimpleNet [21] 0.867 0.574 0.876 0.723 0.766 0.531 0.244 0.801 0.828 0.757 0.737 0.918 0.718
EfficientAD [1] 0.982 0.897 0.888 0.822 0.895 0.968 0.982 0.945 0.948 0.950 0.982 0.982 0.937
RD++ [29] 0.964 0.959 0.699 0.642 0.919 0.977 0.979 0.932 0.938 0.957 0.949 0.967 0.907
FBFT (Ours) 0.979 0.963 0.971 0.908 0.944 0.971 0.961 0.965 0.939 0.910 0.935 0.972 0.952
AUPRO@5% PatchCore [25] 0.823 0.402 0.783 0.491 0.497 0.140 0.133 0.799 0.609 0.399 0.592 0.831 0.542
SimpleNet [21] 0.660 0.396 0.650 0.444 0.397 0.221 0.121 0.611 0.579 0.385 0.446 0.715 0.469
EfficientAD [1] 0.897 0.675 0.715 0.582 0.585 0.839 0.897 0.779 0.775 0.782 0.897 0.897 0.777
RD++ [29] 0.859 0.826 0.505 0.384 0.749 0.872 0.879 0.789 0.811 0.827 0.739 0.851 0.758
FBFT (Ours) 0.881 0.833 0.851 0.674 0.751 0.848 0.839 0.834 0.746 0.684 0.664 0.840 0.787
A.6 Training time.
We provide in Tab. 10 the average time needed per class to train every framework given the number of epochs reported in their official implementations. These timings have been computed using the same hardware employed for all our experiments.
Table 10: Training time required on the VisA dataset. Average training time in hours per class. All methods are trained and tested at high-resolution.
Algorithm PatchCore [25] SimpleNet [21] EfficientAD [1] RD++ [29] FBFT (Ours)
Training time 1.212 6.266 7.783 28.767 2.361
A.7 Implementation employed for the competitors and their licenses
For all the competitors [25, 21, 29], except EfficientAD [1], we employed their official implementations. As far as it concerts EfficientAD, which does not provide an official repository, we leverage on an implementation that obtains the most similar results with respect to the values reported in their manuscript [1]. In particular:
A.8 License for the employed datasets
The VisA dataset [34] is released under the Creative Commons Attribution (CC BY 4.0) license. The MVTec AD dataset [4] is released under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License (CC BY-NC-SA 4.0).
A.9 Additional qualitative results on the VisA and MVTec AD datasets.
As anticipated in the main paper, we show in Fig. 7 some additional qualitative results for the remaining classes of the VisA dataset which have not been reported in Fig. 5. As already highlighted in Sec. 5, the anomaly maps produced by our solution provide a more localized response for the anomalies, compared to EfficientAD [1].
Additionally, in Fig. 8 we show some qualitative examples of the anomaly map produced by our model on the MVTec AD dataset. Also in this scenario, our method provides more localized anomaly scores, motivating the segmentation performance gap in terms of both AUPRO@30% and AUPRO@5%.
macaroni2 pcb1 pcb2 pcb3 pcb4 pipe_fryum
RGB
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EfficientAD-M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FBFT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: VisA dataset qualitative results. All methods are trained and tested at high resolution.
bottle cable capsule carpet grid hazelnut leather
RGB
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EfficientAD-M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FBFT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
metal_nut pill tile toothbrush transistor wood zipper
RGB
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EfficientAD-M
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FBFT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: MVTec AD dataset qualitative results.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes] .
  4. Justification: the abstract and the introduction outline both the benchmark and the method that are presented in detail in the rest of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes] .
  9. Justification: the limitations are briefly discussed in Sec. 6 of the paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A] .
  14. Justification: the paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes] .
  19. Justification: all the implementation details of the method are specified and random seeds are fixed across all experiments. Moreover, the code will be released upon eventual acceptance with both train and evaluation scripts, along with the model checkpoints.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes] .
  24. Justification: All data employed in the experiments are publicly available and licensed. The code will be released upon eventual acceptance with both train and evaluation scripts, along with the model checkpoints. Code for the baseline with an official implementation will not be provided but the relative information and repository are described in the supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes] .
  29. Justification: implementation details are provided in Sec. 5. Data splits are defined by the respective datasets.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No] .
  34. Justification: experiments have not been run multiple times since it would be too time and resource-consuming. The inference is deterministic. Random seeds have been fixed to ensure stability in the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes] .
  39. Justification: training and inference times are reported for all the competitors and our method.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes] .
  44. Justification: we reviewed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] .
  49. Justification: there is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: the paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: the creators and original owners of assets are properly credited and license and terms of use explicitly mentioned in the supplemental material and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code released upon eventual acceptance will provide a proper documentation with instructions to run both train and evaluation scripts, along with the model checkpoints.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] .
  69. Justification: the paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] .
  74. Justification: the paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Figure 8: MVTec AD dataset qualitative results.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes] .
-  4. Justification: the abstract and the introduction outline both the benchmark and the method that are presented in detail in the rest of the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes] .
-  9. Justification: the limitations are briefly discussed in Sec. 6 of the paper.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A] .
-  14. Justification: the paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes] .
-  19.
-Justification: all the implementation details of the method are specified and random seeds are fixed across all experiments.
-Moreover, the code will be released upon eventual acceptance with both train and evaluation scripts, along with the model checkpoints.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes] .
-  24.
-Justification: All data employed in the experiments are publicly available and licensed.
-The code will be released upon eventual acceptance with both train and evaluation scripts, along with the model checkpoints.
-Code for the baseline with an official implementation will not be provided but the relative information and repository are described in the supplemental material.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes] .
-  29. Justification: implementation details are provided in Sec. 5. Data splits are defined by the respective datasets.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No] .
-  34.
-Justification: experiments have not been run multiple times since it would be too time and resource-consuming.
-The inference is deterministic.
-Random seeds have been fixed to ensure stability in the experiments.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes] .
-  39. Justification: training and inference times are reported for all the competitors and our method.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes] .
-  44. Justification: we reviewed the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A] .
-  49. Justification: there is no societal impact of the work performed.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A] .
-  54. Justification: the paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: the creators and original owners of assets are properly credited and license and terms of use explicitly mentioned in the supplemental material and properly respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The code released upon eventual acceptance will provide a proper documentation with instructions to run both train and evaluation scripts, along with the model checkpoints.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A] .
-  69. Justification: the paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A] .
-  74. Justification: the paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.